Given the function f(x) = 8x2 (it’s supposed to be x squared) +5x-5. Calculate the following values:
f(0) =
f(2) =
f(-2) =
f(x+1) =
f(-x) =

Answer by Alan3354(69443)

Given the function f(x) = 8x^2+5x-5. Calculate the following values:
---
Use ^ (Shift 6) for exponents.
---------
f(0) =
f(2) =
f(-2) =
===================================
f(x+1)
Wherever you see x, sub (x+1)
f(x) = 8x^2+5x-5
f(x+1) = 8(x+1)^2 + 5(x+1) - 5
= 8x^2 + 16x + 8 + 5x+5 - 5
= 8x^2 + 21x + 8
------
I did the "hardest one."
Do the rest the same way.
------------
f(-x) =